# VizGroup: An AI-Assisted Event-Driven System for Real-Time Collaborative Programming Learning Analytics

### v1

@@ 5.2.1. Trackers (Figure  5)
5.2.2. Alerts (Figure  6)
Drawing from previous research on the use of visual aids to signal the current status of groups in online breakout rooms and alerts that monitors abnormal activity in complex systems (Sato et al., 2023; Bunch et al., 2004), Alerts function as a tool for tracking when student behaviors or group interactions surpass limits set by instructors. They encapsulate the activity level, group conversation topic, the collective pass rate, team structure for group analysis, and code issues, activity levels, and individual pass rates for individual analysis (Table  1). However, The alert card shows the alert type (group or individual) (Figure 6.2), the reason why the alert is created (Figure  6.3), the current criteria for the alert (Figure  6.4), the list of data points that currently satisfies the criteria (Figure  6.6) and the total number data points in the list (Figure  6.1).Instructors can modify the types of activity attributes they want to be alerted about (Figure  6.10). For each attribute, clicking on the highlighted area reveals a drop down menu (Figures  6.4), allowing instructors to modify its threshold values (Figure  6.5).
Alerts can be activated spatially or temporally. For spatially, the alert is activated when the quantity of students surpassing a predetermined threshold exceeds n (e.g., alert instructors when the number of students that has a pass rate lower than 50% and an activity level from 0 to 3 is more than 10; Figure  8). Imagine a digital map detailing the layout of a convention venue, where each attendee is represented by a red dot. This interactive tool offers organizers a bird’s-eye view of crowd distribution, enabling them to monitor shifts in attendee density across different areas of the venue in real-time. Such a system allows for the immediate identification of overcrowded sections. Upon detecting areas that exceed a predetermined threshold of attendees, the system can automatically alert organizers. This prompts them to take proactive measures, such as deploying additional staff or resources to manage the situation effectively. Similarly, this concept extends to Alerts, which is based on a 5-dimensional model of student activity attributes. Instructors receive notifications when the number of students that falls under the predefined student activity range, allowing for timely interventions.
@@ 5.2.2. Alerts (Figure  6)
For temporally, the alert is activated when specific groups or students surpass the predetermined threshold for a duration exceeding t seconds or minutes (e.g., alert instructors of all students that discuss correcting function logic and syntax errors in their group conversation and have a pass rate between 0% and 41% for over 2 minutes; Figure  7). Consider setting an alarm clock, where an clock will sound when the time is up. Alerts uses a similar idea where each datapoint that meets the user defined criteria has an internal alarm clock, keeping track the amount of time they meet the defined criteria. When the time is up for each individual datapoint, namely the period of time they meet the criteria exeeds the user defined time, it will push an alert which will display the datapoint to the instructor for further investigation. After defining the alert criteria, a curated list of groups and students that meets the threshold enables instructors to examine associated group details (Figure  6.6). Instructors can select Moreover, VizGroup provides an option for instructors to preview and visualize data points that currently align with alert criteria, offering a macroscopic view of the tracked groups’ spatial distribution (Figure  6.8). By clicking on the time or number icons (Figure  6.7), instructors can transition between these alert types.
After confirming the alert thresholds by clicking the green check mark, the alert transitions to an Active Notification (Figure  6.9). The alert then invokes a flashing animation to ensure that instructor is aware of any updates (Figure  6.11).
@@ 5.3.1. User Interaction
VizGroup generates suggested notifications for the scatter plot view currently in use. Specifically, group alerts will be suggested when instructors are in group and structure view and individual alerts will be suggested when instructor is in individual view. Instructors could also create notifications by selecting areas of interest on the scatter plot. When they click on a specific data point, VizGroup derives a suggestion on student activity attributes to be tracked based on the displayed view of the scatter plot. For the Structure View, it would suggest alerts based on the selected group conversation topics and the number of team members active in the group (Figure  9.2). For the Group and Individual view, it would suggest alerts based on group or individual activity level and pass rate. When the instructor highlights an area in the scatter plot, the x and y range of the selected area will be automatically applied towards the suggested notification, enabling instructors to select data points to keep track of through direct manipulation. Instructors can use such suggestions as templates to further modify alerts to suit their needs. Similarly, when instructors inspect the rows of the aggregated topics and errors in the detailed information table, suggestions are based on the expanded row (Figure  9.3).
@@ 5.3.2. Historic Student Activity Data
We asked the LLM to analyze each group’s history of submission attempts, group conversation logs and current status data such as group pass rate, team activity, the summarized group topic, and the number of members participating in the group discussion. After this holistic evaluation of group performance, we asked the LLM to identify specific issues each group was facing. Some examples of issues were no active conversation despite submission attempts or group members asking for help but the conversation lacked problem solving.
@@ 6.1. Participants
We recruited 12 participants (5 females and 7 males) who had experience teaching programming courses at four-year universities via personal networks, local mailing lists, and snowball sampling. During the study, participants were asked to interact with two of the three systems to inspect student behavior that was collected in a large programming course at our institute. Each participant was compensated with \$25 USD for their time and effort.
@@ 6.2.1. Live Simulation
During the study, participants watched live playback of the session to simulate a real-time, in-class peer instruction session. To ensure the data participants interacted with was authentic, we used real data captured from a large-scale introductory level university programming course’s peer instruction session that contained 111 students that were divided into 37 groups.
The grouping strategy was to gather students who passed the test with those who did not and ensure each group had at least one student who passed the test. During the peer instruction process, students accessed the system through their laptops. They were not able to see group members’ real names and their code submissions. The programming exercise for the session was an introductory Python problem to count the number of elements under 100 in a given list of numbers.
@@ 6.2.2. Conditions
  • Baseline (A): a baseline version of VizGroup without any intelligent or automatic features such as notifications, topic summaries, and team activities. The system still contained the interactive visualization.
  • VizGroup without suggested notification (B2): an ablated version of VizGroup without the context-aware suggestion feature. This version still had the notification feature.
  • VizGroup (B1): a full AI-assisted version of VizGroup with all its features.
Figure 11. Survey responses after the open-ended tasks. For the ease of use and usefulness questions, 1 was very negative and 7 was very positive. For the NASA-Task questions  (Hart and Staveland, 1988), 1 indicated very low levels of mental demand, temporal demand, successful performance, effort, and frustration.
@@ 6.2.3. Tasks
The quiz questions were designed to evaluate the system’s ability to enable participants to overview, narrow down, and track nuances throughout the class. The open-ended task sought to evaluate participants’ experiences with the customized usage of the notification feature to track students’ behavior and groups’ interaction over time. These questions were derived from our formative study and iterated on. We counterbalanced the conditions and problem sets to reduce potential learning effect.
@@ 6.3.1. Quantitative Results
We combined the results from the two conditions with the notification features as the quiz questions did involve the use of the historic-based suggestions (Table 2). The results showed that participants who used VizGroup had shorter times, higher recall, and higher precision on Q1, and significantly shorter time and higher recall on Q2.
For the open-ended task, participants created more notifications for both students (VizGroup: $Mean=2.67,\sigma=1.37$, Baseline: $Mean=2.00,\sigma=1.26$) and groups (VizGroup: $Mean=3.00,\sigma=2.10$, Baseline: $Mean=2.50,\sigma=1.64$). There was not a statistically significant difference in the number of created notifications.
@@ 6.3.3. Context-Aware Suggestions Increased the Diversity of Notifications
During the open-ended tasks, the notifications created with VizGroup were more diverse than the ablated version of VizGroup (Number of notifications containing different types of criteria–VizGroup: $Mean=4.50,\sigma=2.66$, VizGroup without suggestion: $Mean=3.33,\sigma=1.63$; not statistically significant). For instance, before inspecting the suggested notification, P3 only created a notification about Pass Rate and Activity Level, however, after inspecting and adapting suggested notifications, P3’s notifications covered all of the criteria.
@@ 6.3.4. VizGroup Aided Instructors In Identifying and Tracking Patterns and Work in Parallel
Nine participants mentioned the notification (with or without suggestion) would be helpful to ‘‘teach a 100/200 people programming class”. “(VizGroup) Makes it much more convenient and feasible so you don’t have to keep checking.”(P10).
@@ 6.3.5. VizGroup Enables Instructors to be Prepared Before Sending Support
Participants (P9, P2, P3, P6) also found that based on the meaning of the constraints applied to the alert, they could make more sense of how to help students with different types of difficulties. P6 suggested that “The topics filter can help professors/TAs prepare beforehand to help students instead of just going up to them and asking what they’re struggling on.”, while P9 also expressed “having different alerts to monitor different groups of students’ behavior, so I can tell TA who can offer help in those different situation.”
@@ 7.4. Limitations
There are several limitations to our user study. First, although we used authentic code and peer discussion data, we conducted the study in a simulated setting rather than in a real living classroom, which reduced the psychological intensity and potential for distraction among participants. Second, the discussion data is only captured from one session during one class, which means that the result may not be generalizable to other contexts. Additionally, there are some limitations with our system. First, the choice of collaborative learning aspects used for generating notifications was limited to Pass Rate, Activity Level, Code Issue, Conversation Topic, and Team Structure. Second, it is difficult to scale the number of notifications to help users understand and navigate the system.
@@ 8. Conclusion
In this paper, we introduced VizGroup, a novel system that helps programming instructors create notifications, such as alerts and trackers, to better manage students’ in-class Peer Instruction activity. It achieves this by leveraging LLMs to adapt time-sensitive contextual information, such as user interactions and historical data changes, and recommends urgent notification units for instructors to monitor. Through our comparison study, we found that with our notification recommendations, participants discovered patterns that they were previously unaware of, and that the created notifications covered a more diverse range of collaborative learning metrics that led to changes in their creation strategies. Our work contributes new understandings and design lessons to real-time, large-scale collaborative learning analytics for domain experts who are novice users of Visual Analytics. Collectively, this opens a new door for building systems that support in-class collaborative learning at scale.

### v2

@@ 5.2.1. Trackers (Figure  6)
5.2.2. Alerts (Figure  7)
Drawing from previous research on the use of visual aids to signal the current status of groups in online breakout rooms and alerts that monitors abnormal activity in complex systems (Sato et al., 2023; Bunch et al., 2004), Alerts function as a tool for tracking when student behaviors or group interactions surpass limits set by instructors. They encapsulate the activity level, group conversation topic, the collective pass rate, team structure for group analysis, and code issues, activity levels, and individual pass rates for individual analysis (Table  1). However, the alert card shows the alert type (group or individual) (Figure 7.2), the reason why the alert is created (Figure  7.3), the current criteria for the alert (Figure  7.4), the list of data points that currently satisfies the criteria (Figure  7.6) and the total number data points in the list (Figure  7.1). Instructors can modify the types of activity attributes they want to be alerted about (Figure  7.10). For each attribute, clicking on the highlighted area reveals a drop down menu (Figures  7.4), allowing instructors to modify its threshold values (Figure  7.5).
Alerts can be activated spatially or temporally. For spatially, the alert is activated when the quantity of students surpassing a predetermined threshold exceeds n (e.g., alert instructors when the number of students that has a pass rate lower than 50% and an activity level from 0 to 3 is more than 10; Figure  9).
These alerts are based on a 5-dimensional model of student activity attributes. Instructors receive notifications when the number of students that falls under the predefined student activity range, allowing for timely interventions.
@@ 5.2.2. Alerts (Figure  7)
Temporally alerts are activated when specific groups or students surpass the predetermined threshold for a duration exceeding t seconds or minutes (e.g., alert instructors of all students that discuss correcting function logic and syntax errors in their group conversation and have a pass rate between 0% and 41% for over 2 minutes; Figure  8). Consider setting an alarm clock, where a clock will sound when the time is up. Alerts use a similar idea where each datapoint that meets the user defined criteria has an internal alarm clock, keeping track the amount of time they meet the defined criteria. When the time is up for each individual datapoint, namely the period of time they meet the criteria exeeds the user defined time, it will push an alert which will display the datapoint to the instructor for further investigation. After defining the alert criteria, a curated list of groups and students that meets the threshold enables instructors to examine associated group details (Figure  7.6). Moreover, VizGroup provides an option for instructors to preview and visualize data points that currently align with alert criteria, offering a macroscopic view of the tracked groups’ spatial distribution (Figure  7.8). By clicking on the time or number icons (Figure  7.7), instructors can transition between these alert types.
After confirming the alert thresholds by clicking the green check mark, the alert transitions to an Active Notification (Figure  7.9). The alert then invokes a flashing animation to ensure that instructor is aware of any updates (Figure  7.11).
@@ 5.3.1. User Interaction
VizGroup generates suggested notifications for the scatter plot view currently in use. Specifically, group alerts will be suggested when instructors are in group and structure view and individual alerts will be suggested when instructor is in individual view. Instructors could also create notifications by selecting areas of interest on the scatter plot. When they click on a specific data point, VizGroup derives a suggestion on student activity attributes to be tracked based on the displayed view of the scatter plot. For the Structure View, it would suggest alerts based on the selected group conversation topics and the number of team members active in the group (Figure  10.2). For the Group and Individual view, it would suggest alerts based on group or individual activity level and pass rate. When the instructor highlights an area in the scatter plot, the x and y range of the selected area will be automatically applied towards the suggested notification, enabling instructors to select data points to keep track of through direct manipulation. Instructors can use such suggestions as templates to further modify alerts to suit their needs. Similarly, when instructors inspect the rows of the aggregated topics and errors in the detailed information table, suggestions are based on the expanded row (Figure  10.3).
@@ 5.3.2. Historic Student Activity Data
Following prior work on using LLM to identify challenges in students’ learning and collaborations (Suraworachet et al., 2024), we asked the LLM to analyze each group’s history of submission attempts, group conversation logs and current status data such as group pass rate, team activity, the summarized group topic, and the number of members participating in the group discussion. After this holistic evaluation of group performance, we asked the LLM to identify specific issues each group was facing. Some examples of issues were no active conversation despite submission attempts or group members asking for help but the conversation lacked problem solving.
@@
The primary objective of utilizing LLMs in the suggestion generation process is to identify complex patterns in students’ code and text-based discussions, transforming these into easily inspectable suggestions to support instructors’ learning analytics. To enhance accuracy, we also adopted state-of-the-art prompting engineering techniques such as few-shot prompts (Brown et al., 2020) and AI-chains (Wu et al., 2022).
@@ 6.1. Participants
We recruited 12 participants (5 females and 7 males) who had experience teaching programming courses at four-year universities via personal networks, local mailing lists, and snowball sampling. During the study, participants were asked to interact with VizGroup under different conditions to inspect student behavior that was collected in a large programming course at our institute. Each participant was compensated with \$25 USD for their time and effort.
@@ 6.2.1. Live Simulation
During the study, participants watched live playback of the session to simulate a real-time, in-class peer instruction session. To ensure the data participants interacted with was authentic, we used real data captured from a large-scale introductory level university programming course’s peer instruction session that contained 111 students that were divided into 37 groups. The grouping strategy was to gather students who passed the test with those who did not and ensure each group had at least one student who passed the test. During the peer instruction process, students accessed the system through their laptops. They were not able to see group members’ real names and their code submissions. The programming exercise for the session was an introductory Python problem to count the number of elements under 100 in a given list of numbers.
@@ 6.2.2. Conditions
  • Baseline (A): a baseline version of VizGroup without any intelligent or automatic features such as notifications, topic summaries, and team activities (i.e., without (2), (3), and topics in (5) and (7) in Figure 3. The system still contained the interactive visualization.
  • VizGroup without suggested notification (B2): an ablated version of VizGroup without the context-aware suggestion feature. This version still had the notification feature, but all notifications needed to be manually created.
  • VizGroup (B1): a full AI-assisted version of VizGroup with all its features. Notification suggestions are dynamically generated in B1. Interaction-based suggestions were triggered by users’ interaction, while the system displayed a new historic-based suggestion every 15 seconds.
Figure 12. Survey responses after the open-ended tasks. For the NASA-Task questions  (Hart and Staveland, 1988) (up), 1 indicated very low levels of mental demand, temporal demand, successful performance, effort, and frustration. For the ease of use and usefulness questions (down), 1 was very negative and 7 was very positive.
@@ 6.2.3. Tasks
The quiz questions were designed to evaluate the system’s ability to enable participants to overview, narrow down, and track nuances throughout the class. The open-ended task sought to evaluate participants’ experiences with the customized usage of the notification feature to track students’ behavior and groups’ interaction over time. These questions were derived from our formative study and iterated on. While participants watched the playbacks from one session under both conditions, we used the same clips from the playback for open-ended tasks (between-subject) and used different clips with counterbalanced order for the quiz questions (within-subject) to reduce overhead costs (e.g., getting familiar with an exercise). We also counterbalanced the conditions to reduce potential learning effect.
@@ 6.3.1. Quantitative Results
We combined the results from the two conditions (B1, B2) with the notification features as the quiz questions did not involve the use of the historic-based suggestions (Table 2). The results showed that participants who used VizGroup had shorter times, higher recall, and higher precision on Q1, and significantly shorter time and higher recall on Q2.
In the open-ended tasks, participants using VizGroup received 18.26 ($\sigma=1.21$) notification suggestions on average. Among all the suggestions, there were 13 types of notifications categorized by activity attributes while 11 of them were accepted by participants. Participants created more notifications for both students (VizGroup: $Mean=2.67,\sigma=1.37$, VizGroup without suggestion: $Mean=2.00,\sigma=1.26$) and groups (VizGroup: $Mean=3.00,\sigma=2.10$, VizGroup without suggestion: $Mean=2.50,\sigma=1.64$). In total, there were $25.93\%$ more notifications created using VizGroup (VizGroup: $Mean=5.67,\sigma=3.44$, VizGroup without suggestion: $Mean=4.50,\sigma=2.81$). There was not a statistically significant difference in the number of created notifications. Of 5.67 created notifications using VizGroup, 2.50 ($\sigma=2.95$) of them are from historic-based suggestions, 2.67 ($\sigma=1.21$) of them are from interaction-based suggestions, and 0.50 ($\sigma=1.22$) are manually created.
@@ 6.3.3. Context-Aware Suggestions Increased the Diversity of Notifications
During the open-ended tasks, the notifications created with VizGroup were $35.14\%$ more diverse than the ablated version of VizGroup (Number of notifications containing different types of criteria–VizGroup: $Mean=4.50,\sigma=2.66$, VizGroup without suggestion: $Mean=3.33,\sigma=1.63$; not statistically significant). For instance, before inspecting the suggested notification, P3 only created a notification about Pass Rate and Activity Level, however, after inspecting and adapting suggested notifications, P3’s notifications covered all of the criteria.
@@ 6.3.4. VizGroup Aided Instructors In Identifying and Tracking Patterns and Work in Parallel
Nine participants mentioned the notification (with or without suggestion) would be helpful to “teach a 100/200 people programming class”. “(VizGroup) Makes it much more convenient and feasible so you don’t have to keep checking.” (P10).
@@ 6.3.5. VizGroup Enables Instructors to be Prepared Before Sending Support
Participants (P2, P3, P6, P9) also found that based on the meaning of the constraints applied to the alert, they could make more sense of how to help students with different types of difficulties. P6 suggested that “The topics filter can help professors/TAs prepare beforehand to help students instead of just going up to them and asking what they’re struggling on.”, while P9 also expressed “having different alerts to monitor different groups of students’ behavior, so I can tell TA who can offer help in those different situation.”
@@ 7.4. Limitations
There are several limitations to our user study. While we used authentic code and peer discussion data, we conducted the study in a simulated setting rather than in a real living classroom, which reduced the psychological intensity and potential for distraction among participants. Despite a high level of participants-reported usefulness and increases in both quantity and diversity of notifications using VizGroup in our lab study, instructors did not validate/interact with real students. Future research can also investigate instructors’ processes for validating AI-generated content, particularly student-related information, in real-time classroom settings.
The discussion and submission data is only captured from one session during one class, which means that the result may not be generalizable to other contexts. Likewise, our evaluation of VizGroup used a dataset from a large programming lecture, limiting the scope to large-scale data. We suspect that participants spent less time on pattern identification and verifying notifications when dealing with smaller datasets, as there was less information to process. Future research can further explore the relationship between data size and visualization complexity.
Additionally, there are some limitations to our system. First, the choice of collaborative learning aspects used for generating notifications was limited to Pass Rate, Activity Level, Code Issue, Conversation Topic, and Team Structure. Second, it is difficult to scale the number of notifications to help users understand and navigate the system. Moreover, the use of LLM could raise several concerns regarding stability, transparency, and trust. Though our approach adopted prior work’s methods (Dai et al., 2023; Suraworachet et al., 2024) that was examined through metrics (e.g., F1, NDCG@K), further evaluation of LLMs’ capabilities in learning analytics is necessary. Lastly, our approach to generating notification suggestions faces limitations in handling context length when there are numerous messages or submissions. This issue could be mitigated with context management techniques in future work.
@@ 8. Conclusion
In this paper, we introduced VizGroup, a novel system that helps programming instructors create notifications, such as alerts and trackers, to better manage students’ in-class Peer Instruction activity. It achieves this by leveraging LLMs to adapt time-sensitive contextual information, such as historical data changes, and recommends urgent notification units for instructors to monitor. Through our comparison study, we found that with our notification recommendations, participants discovered patterns that they were previously unaware of, and that the created notifications covered a more diverse range of collaborative learning metrics that led to changes in their creation strategies. Our work contributes new understandings and design lessons to real-time, large-scale collaborative learning analytics for domain experts who are novice users of Visual Analytics. Collectively, this opens a new door for building systems that support in-class collaborative learning at scale.